Is the following trinomial a perfect square?
x^2- 14x +49
I came up with the answer of yes. Is this correct? Thanks.

Answer by wodkit02(9)

To see if the equation is factorable, I would use the following method:
ax^2 + bx + c = 0
Multiply a and c and figure out what two factors of ac add up to b.
Our problem: x^2- 14x +49
a * c = 1 * 49 = 49
factors of 49 are 1 x 49, -1 x -49, -7 x -7, and 7 x 7
Since -7 + -7 = -14, and b = -14, we get the factors (x-7)(x-7).
Because (x-7)(x-7) = (x-7)^2, this trinomial is a perfect square.